OpenStudy (anonymous):

the length of a rectangular is twice the width. The area is 128yd^2. Find the length and width.

OpenStudy (btaylor):

A = L x W L = 2W A = 2W x W 128 = 2W^2 64 = W^2 W = 8 --> L=16